A man is 2 years older than his sister. Thirty years ago he was 3 times as old as her. How old is the man now?
Answer by ankor@dixie-net.com(22740)

A man is 2 years older than his sister.
Thirty years ago he was 3 times as old as her.
How old is the man now?
;
let m = the man's age
let s = the sisters' age
:
Write an equation for each statement
:
"A man is 2 years older than his sister.'
m = s + 2
:
" Thirty years ago he was 3 times as old as her."
m - 30 = 3(s - 30)
m - 30 = 3s - 90
m = 3s - 90 + 30
m = 3s - 60
Replace m with (s+2), from the 1st statement
s + 2 = 3s - 60
2 + 60 = 3s - s
62 = 2s
s = 62/2
s = 31 yrs is the sister's age
then
m = 31 + 2
m = 33 is the man's age
:
:
Confirm this using the statement:
"Thirty years ago he was 3 times as old as her."
33 - 30 = 3(31 - 30)
3 = 3(1)